OpenStudy (anonymous):

The perimeter of an isosceles triangle is 29 feet. If the base measures 15 feet what is the measure of the other two sides? Answers: 14 ft 22 ft 11 ft 7 ft

OpenStudy (aravindg):

2x+15=29 2x=29-15=14 x=7